# Modification Takes Courage: Seamless Image Stitching via Reference-Driven Inpainting

### v1

@@ 3.1 Problem Definition
$I_{wr},I_{wt}=\mathcal{W}(I_{r},\mathcal{H}),\mathcal{W}(I_{r},\mathbb{I}),$ (1)
@@ 4.1 Experimental Setup
Baseline method. Selecting appropriate baseline methods is a little difficult, because our method is the first to propose a fusion and rectangling unified model that can be trained on the unsupervised image stitching dataset, a groundbreaking advancement in the image stitching. As mentioned in Sec.2.1, UDIS [22] and UDIS++ [25] can only perform up to the fusion stage and require integration with another rectangling model to complete the entire image stitching pipeline. However, the SOTA rectangling models [23, 36, 38, 46] are all supervised and can only be trained on a task-specific dataset DIR-D [23], rather than the real-world image stitching dataset UDIS-D. This limitation makes it difficult to compare these methods with RDIStitcher under relatively fair conditions. To address this issue, we propose a variant of RDIStitcher, called RDIStitcher-R, which works only as a rectangling model trained with the same settings as RDIStitcher. With this, we can combine the RDIStitcher-R with UDIS and UDIS++ to establish baselines. In summary, the baselines include the SOTA recon-based method UDIS with RDIStitcher-R (UDIS+R) and the SOTA seam-based method UDIS++ with RDIStitcher-R (UDISplus+R). The implementation details of RDIStitcher-R can be found in the Supplementary Material.
In addition, the baselines also include the SOTA inpainting-based method SRStitcher [48]. Although the SRStitcher method does not involve any training or fine-tuning, which makes the comparison somewhat less fair, we believe it is necessary to include the results of this most important related work. The experimental results demonstrate that our method significantly outperforms SRStitcher. This result represents a noteworthy advancement in the inpainting-based image stitching framework.
Finally, we also implement our self-supervised training method on LeftRefill [6] as a baseline.
@@ Conclusion
Other metrics. In addition to the MLLMs-based image quality metrics, we also introduce the Content Consistency Score (CCS) from SRStitcher [48] and the real user evaluation from UDIS [22] with the same settings.
@@ 4.3 Quantitative Evaluation
Single-image evaluation. Table.3 presents the comparative experimental results of SIQS-Q, SIQS-G, and CCS on the UDIS-D test set with 1,106 sample pairs. Our method surpasses other methods in terms of both quality and content consistency.
@@ 4.4 Qualitative Evaluation
4.5 Discussion
Limitation of LeftRefill. The results presented in Table.3 and Fig.5 indicate that LeftRefill [6] performs significantly worse than other methods. By observing and analyzing the experimental results, we believe that LeftRefill struggles to accurately differentiate between fusion and rectangling tasks. As a result, it inpaints the coarse rectangling image into the stitched image, leading to significant errors. While LeftRefill achieves reference-driven inpainting with minimal parameter adjustments (hundreds of kilobytes), image stitching is a complex mixed task that includes both fusion and rectangling. This complexity poses challenges for LeftRefill. Therefore, in developing RDIStitcher, we use more fine-tuned parameters (dozens of megabytes) and a different fine-tuned network structure than LeftRefill. Although the saved LoRA weights are slightly larger, we believe that this difference does not significantly affect the usability of our method. In addition, despite the increased number of adjusted parameters, the memory required for training our method is much lower than that for LeftRefill, which may be caused by some operations required by LeftRefill on the attention modules.
5 Conclusion and Future Work
In this paper, we propose an image stitching model based on reference-driven inpainting, which achieves remarkable seamless stitching results on challenging scenes. Extensive experiments demonstrate the effectiveness of our model, particularly highlighting its strong generalization performance in zero-shot scenarios. However, through a comprehensive analysis of the experimental results, we believe that there are three critical open issues to be addressed in the future:
(1) Stability of content generated in rectangling regions. Table.3 shows that the standard deviation of content consistency in our method is relatively higher compared to that of SRStitcher. Because our method tries to provide specific images instead of local blur, which results in reduced stability. In the future, it may be possible to consider using a different loss to optimize the stability of rectangling regions.
(2) Accuracy of MLLMs-based stitched image quality metrics. Although the MLLMs-based metrics provide new insights into stitched image metrics, they still have limitations (see the Supplementary Material for details). However, we believe that MLLMs-based measurement is a promising research direction, and we are considering more further in-depth studies to introduce more comprehensive MLLMs and to develop prompts for comparison.
(3) Reliability of the generation-based methods. Some researchers express concern about the reliability of applying generation-based models to traditional vision problems, likening them to a lottery machine. However, we argue that, compared to previous methods, our method at least confirms that a big prize exists in the solution space. The future challenge is finding a more precise way to locate this big prize.
@@ Appendix A Overview
  • Source code for our work, including train, inference, and MLLMs-based metrics. Please see the code document;
@@
  • More Results of Qualitative Evaluations (Sec.D);
@@
  • More Details of MLLMs-based Metrics (Sec.F).
@@ C.2 Ablation Study of RDIStitcher-R
The Table.4 shows the ablations of USID+RDIStitcher-R (UDIS+R) based on the content consistency score CCS [48] on the UDIS-D test set. The comparisons include UDIS+stable diffusion inpainting model [2] without fine-tuning (UDIS+SD2) and UDIS+RDIStitcher-R without coarse rectangling (UDIS+RWCR). The results demonstrate that RDIStitcher-R effectively reduces the likelihood of abnormal content.
@@
Appendix D More Results of Qualitative Evaluations
Due to space limitations, we only present a limited number of qualitative evaluation results in the main paper. Here, to further examine the generalization capabilities of the RDIStitcher, We provide additional qualitative results obtained on the APAPdataset [40], REWdataset [15], and SPWdataset [18].
The additional qualitative results on the APAPdataset [40] are presented in Fig.8. Note that in the fifth case, the stitched image produced by our method appears to exhibit some errors in the rendering of Chinese characters. We speculate that this is likely due to the fact that the underlying T2I model employed is not optimized for Chinese language processing. As a result, the model may not be well-equipped to handle the accurate representation of Chinese characters. This paper utilizes the Stable Diffusion 2 inpainting version, acknowledging that Stable Diffusion 3 has significantly improved character generation. However, since no open-source inpainting version is currently available for Stable Diffusion 3, we are unable to test our method on it. Based on research trends, we believe that the character generation issue will not remain a major bottleneck and will be effectively resolved soon.
Moreover, the additional qualitative results on the REWdataset [15] are shown in Fig.9. And the additional qualitative results on the SPWdataset [18] are shown in Fig.10. The scenarios of SPWdataset are a relatively simple, and it is evident that most methods demonstrate good performance.
@@ Appendix E More Information of Ablation Studies
E.1 Different Hyper-parameter Configurations
LoRA Configurations LoRA [13] is considered one of the most influential technologies in recent years, with numerous studies exploring its mechanisms. Due to the limited space of the main paper, all parameters of LoRA are not introduced in detail. Here, we provide additional parameters: rank 8, alpha constant 16, dropout rate 0.1, and bias none. For more details, please refer to the source code we have provided.
We conducted an ablation test by varying the LoRA rank while keeping other parameters fixed. The results are shown in the Table.5. However, we quickly realized that this isolated exploration does not definitively demonstrate that a LoRA rank 8 represents the optimal parameter setting for our specific task. A comprehensive exploration of LoRA cnfigurations is complex and requires significant hardware resources, which is beyond the scope of this study. Therefore, we provide a reference set of LORA parameter values in the main paper, but have not pursued any further ablation analysis regarding the LORA settings.
@@ E.1 Different Hyper-parameter Configurations
Seed robustness We present the qualitative evaluation results with different random seeds in Fig.7. In cases of small parallax, the stitched image produced by our method maintains high consistency. For large parallax, differences in the fusion region are insignificant, and slight changes occur in the rectangling region. Importantly, our method does not introduce significant abnormal content, unlike the original T2I model. In zero-shot tests, the stitched images of uneven hue and small parallax scenes also maintain good consistency, with only minor changes in the rectangling region. In high challenge scenarios, while content changes in the stitched images are more pronounced, our method shows clear advantages in brightness smoothing and content connection compared to previous methods (see Fig.5 of main paper for comparison).
@@ F.3 Experiment Settings of MLLMs-based Metrics
F.4 Limitations of MLLMs-based Metrics
@@
Score mismatch In a well-known case testing LLMs with 9.11 or 9.3 is bigger, most LLMs failed. In our experiment, we also observe a score mismatch phenomenon in MLLMs, where the sum of scores in each aspect can not equal the total score.
@@
Unstable result Our empirical analysis reveals that when the same image is evaluated multiple times by MLLMs, the scores exhibit greater instability, especially GLM-4V-Plus [10]. In some cases, the scores can fluctuate with a variance of up to 4 points.
It is important to acknowledge that MLLMs represent a relatively nascent technology, with the existing public models having been available for testing only in the past two years. This limited timeframe has inevitably led to apparent defects and shortcomings in the current MLLMs. However, it is equally clear that both the research focus and the prevailing research trends strongly indicate the immense potential of MLLMs. Therefore, we believe that the use of MLLMs for stitched image quality evaluation represents a highly promising research direction, and its feasibility should not be hastily dismissed due to the defects present in the current MLLMs.
Figure 8: Additional qualitative results on the APAPdataset[40].
Figure 9: Additional qualitative results on the REWdataset [15].
Figure 10: Additional qualitative results on the SPWdataset [18].

### v2

@@ 3.1 Problem Definition
$I_{wr},I_{wt}=\mathcal{W}(I_{r},\mathcal{H}),\mathcal{W}(I_{t},\mathbb{I}),$ (1)
@@ 4.1 Experimental Setup
Baseline method. Selecting appropriate baseline methods is a little difficult, because our method is the first to propose a fusion and rectangling unified model that can be trained on the unsupervised image stitching dataset, a groundbreaking advancement in the image stitching. As mentioned in Sec.2.1, UDIS [23] and UDIS++ [26] can only perform up to the fusion stage and require integration with another rectangling model to complete the entire image stitching pipeline. However, the SOTA rectangling models [24, 37, 40, 49] are all supervised and can only be trained on a task-specific dataset DIR-D [24], rather than the real-world image stitching dataset UDIS-D. This limitation makes it difficult to compare these methods with RDIStitcher under relatively fair conditions. To address this issue, we propose a variant of RDIStitcher, called RDIStitcher-R, which works only as a rectangling model trained with the same settings as RDIStitcher. With this, we can combine the RDIStitcher-R with UDIS and UDIS++ to establish baselines. In summary, the baselines include the SOTA recon-based method UDIS with RDIStitcher-R (UDIS+R) and the SOTA seam-based method UDIS++ with RDIStitcher-R (UDISplus+R). The test results of RDIStitcher-R on DIR-D are SSIM 0.763 and PSNR 23.56, positioning its performance between the DeepRectangling [24] and Recdiffusion [49]. Implementation details and validation experiments for RDIStitcher-R can be found in the Supplementary Material.
In addition, the baselines also include the SOTA inpainting-based method SRStitcher [51]. Although the SRStitcher method does not involve any training or fine-tuning, which makes the comparison somewhat less fair, we believe it is necessary to include the results of this most important related work. And, we also implement our self-supervised training method on LeftRefill [6] as a baseline.
@@ 4.3 Quantitative Evaluation
In addition to the MLLMs-based image quality metrics, we also introduce the Content Consistency Score (CCS) from SRStitcher [51] and the real user evaluation from UDIS [23] with the same settings in quantitative evaluations.
[…]
@@ 4.4 Qualitative Evaluation
4.5 Consistency Evaluation
To assess whether our method introduces content that deviates from the original images, we use Peak Signal-to-Noise Ratio (PSNR)$\uparrow$ and Structural Similarity Index (SSIM)$\uparrow$ as metrics to compare the consistency of the stitched images’ fusion regions between our method and UDIS. Additionally, the comparison results between UDIS and UDIS++ are provided as references, as shown in the Table.4. The results demonstrate that our method achieves higher consistency in small parallax scenes $D_{S}$, large parallax scenes $D_{L}$, and the entire UDIS-D test set $D_{F}$, confirming its effectiveness in preserving content consistency. Detailed information on $D_{S}$ and $D_{L}$ and further discussion can be found in the Supplementary Material.
Table 4: Consistency evaluation results. We use five seeds and take the mean. The best performer is highlighted by red.
Method $D_{S}$ $D_{L}$ $D_{F}$
UDIS++ vs. UDIS PSNR 32.49 30.71 31.02
SSIM 0.747 0.531 0.602
Ours vs. UDIS PSNR 32.56 31.42 31.66
SSIM 0.824 0.736 0.761
4.6 MLLMs-based Metric Accuracy Evaluation
Referring to work [39], we introduce Spearman’s rank correlation coefficient (SRCC)$\uparrow$ and Pearson linear correlation coefficient (PLCC)$\uparrow$ to evaluate the accuracy of the MLLMs-based metrics and compare them with the advanced NR-IQA methods including Q-align [38], Topiq [7] and UNIQUE (UNIQ) [46] on a specially crafted stitched image quality dataset $D_{quality}$. The results are shown in the Table.5,which indicate that our proposed metrics are more in line with human visual perception. For more details about this experiment and $D_{quality}$, please refer to the Supplementary Material.
Table 5: MLLMs-based metric accuracy evaluation results. The best performer is highlighted by red.
Metric SIQS-Q SIQS-G Q-align Topiq UNIQ
SRCC 0.728 0.634 0.455 0.342 0.299
PLCC 0.685 0.525 0.412 0.335 0.242
5 Conclusion
In this paper, we propose an image stitching model based on reference-driven inpainting, which achieves remarkable seamless stitching results on challenging scenes. In addition, we propose the MLLMs-based metrics for stitched images. Extensive experiments demonstrate the effectiveness of our method and the proposed metrics, and our method highlights its strong generalization performance in the zero-shot scenario.
@@ Appendix A Overview
  • Source code for our method, including train, inference, and MLLMs-based metrics. Please see the code document;
@@
  • More Details of RDIStitcher-R (Sec.C);
  • More Information of Consistency Evaluations (Sec.D);
@@
  • More Details of MLLMs-based Metrics (Sec.F);
  • More Information of Metric Evaluation (Sec.G);
  • More Results of Qualitative Evaluations (Sec.H).
@@ C.2 Ablation Study of RDIStitcher-R
Table.6 shows the ablations of USID+RDIStitcher-R (UDIS+R) based on the content consistency score CCS [51] on the UDIS-D test set. The comparisons include UDIS+stable diffusion inpainting model [2] without fine-tuning (UDIS+SD2) and UDIS+RDIStitcher-R without coarse rectangling (UDIS+RWCR). The results demonstrate that RDIStitcher-R effectively reduces the likelihood of abnormal content.
@@
C.3 Performance of RDIStitcher-R
RDIStitcher-R achieves an SSIM score of 0.763 and a PSNR of 23.56 on the DIR-D dataset, indicating that its performance is situated between the DeepRectangling [24] and Recdiffusion [49]. However, RDIStitcher-R holds a significant advantage in its capability to be trained on unlabeled datasets, making it more versatile in scenarios where labeled data are scarce or difficult to obtain. Additionally, its hardware requirements are notably lower compared to those of Recdiffusion. The performance gap between RDIStitcher-R and Recdiffusion is acceptable in view of these advantages.
Appendix D More Information of Consistency Evaluations
D.1 Information of $D_{S}$ and $D_{L}$
To check the consistency, we manually made the small parallax scenes dataset $D_{S}$ and the large parallax scenes dataset $D_{L}$, both of which are sub-datasets contains 100 samples of the UDIS-D [23], whose details are shown in datasets/ConsistencyDataset.csv in Supplementary Material, where the names of the sample are consistent with those in UDIS-D.
D.2 Experimental results based on masked SSIM and masked PSNR
In addition to the consistency experiments in the main paper, we observed that the parallax regions of the UDIS results are very prone to artifacts leading to low confidence in this region. Therefore, we conduct further tests focusing on the consistency of regions excluding the parallax regions. Specifically, we use the masked Structural Similarity Index (mSSIM) and masked Peak Signal-to-Noise Ratio (mPSNR) as metrics to evaluate image coherence after masking out the parallax regions. The mSSIM and mPSNR were calculated as follows:
$mSSIM$ $=SSIM(I_{s}^{1}\odot M_{o},I_{s}^{2}\odot M_{o}),$ (16)
$mPSNR$ $=PSNR(I_{s}^{1}\odot M_{o},I_{s}^{2}\odot M_{o}),$ (17)
where, $I_{s}^{1}$ and $I_{s}^{2}$ are two images stitched by different methods, such as UDIS and our method. And, $M_{o}$ is the overlap mask, that is $M_{o}=1-(M_{wr}\&M_{wt})$, where &\&& is bitwise and operation.
The results of the consistency experiments based on mSSIM and mPSNR are shown in the Table.7. Note that we don not test the results on $D_{S}$, because in the scenarios with small parallax, the masked area of $M_{o}$ becomes significantly large, leading to most areas of the image without content, thus losing the significance of comparison experiments on the $D_{S}$ dataset.
Table 7: Consistency evaluation results by mSSIM and mPSNR. We use five seeds and take the mean. The best performer is highlighted by red.
Method $D_{L}$ $D_{F}$
UDIS++ vs. UDIS mPSNR 32.90 35.91
mSSIM 0.701 0.838
Ours vs. UDIS mPSNR 33.19 36.04
mSSIM 0.812 0.881
@@ Appendix E More Information of Ablation Studies
E.1 LoRA Configurations
LoRA [14] is considered one of the most influential technologies in recent years, with numerous studies exploring its mechanisms. Due to the limited space of the main paper, all parameters of LoRA are not introduced in detail. Here, we provide additional parameters: rank 8, alpha constant 16, dropout rate 0.1, and bias none. For more details, please refer to the source code we have provided.
We conducted an ablation test by varying the LoRA rank while keeping other parameters fixed. The results are shown in the Table.8. However, we quickly realized that this isolated exploration does not definitively demonstrate that a LoRA rank 8 represents the optimal parameter setting for our specific task. A comprehensive exploration of LoRA cnfiguration is complex and requires significant hardware resources, which is beyond the scope of this study. Therefore, we provide a reference set of LoRA parameter values in the main paper, but have not pursued any further ablation analysis regarding the LoRA settings.
@@ E.1 LoRA Configurations
E.2 Seed robustness
We present the qualitative evaluation results with different random seeds in Fig.7. In cases of small parallax, the stitched image produced by our method maintains high consistency. For large parallax, differences in the fusion region are insignificant, and slight changes occur in the rectangling region. Importantly, our method does not introduce significant abnormal content, unlike the original T2I model. In zero-shot tests, the stitched images of uneven hue and small parallax scenes also maintain good consistency, with only minor changes in the rectangling region. In high challenge scenarios, while content changes in the stitched images are more pronounced, our method shows clear advantages in brightness smoothing and content connection compared to previous methods (see Fig.5 of main paper for comparison).
@@ F.3 Experiment Settings of MLLMs-based Metrics
F.4 The cost of MLLMs-based Metrics
All MLLMs used in this paper offer free access through their APIs, with operating costs below \$0.02 per thousand tokens for both input and output. For comparison, if we were to manually evaluate the UDIS-D test set (1,106 samples), it would take as fast as an hour and a half to perform an evaluation assuming that a stitched images is evaluated for 5 seconds (the actual evaluation time would be longer because the human eye needs a rest period). Therefore, the use of MLLM-based metrics proves to be not only more economical, but also significantly faster than manual inspection.
F.5 Limitations of MLLMs-based Metrics
@@
Appendix G More Information of Metric Evaluation
G.1 Information of $D_{quality}$
We hand-crafted a stitched image dataset for measuring image quality, named $D_{quality}$, containing 50 different stitched images, whose details can be found in datasets/QualityDataset.zip. To enrich the diversity of $D_{quality}$, we intentionally incorporated examples where various stitching methods introduced errors, including severe artifacts, distortions, structural deformations, content anomalies, and noticeable seams. In addition, the dataset includes a selection of well-stitched images to evaluate the ability of the metric to accurately identify high quality stitching results.
G.2 More Details of MLLMs-based Metric Accuracy Evaluation
Since the image stitching datasets are unlabeled, traditional reference-based metrics for assessing image quality cannot be applied. To address this challenge, we adopt a methodology inspired by other image quality evaluation studies, establishing an evaluation benchmark based on mean user scores. Specifically, we invite four volunteers to score the samples in $D_{quality}$ according to the criteria showed in Table.10. Then, we calculate the SRCC and PLCC between the mean user scores and comparison metrics. These correlation coefficients serve as a measure to evaluate the accuracy and reliability of each metric in reflecting the stitched image quality, ensuring its alignment with human perceptual judgement.
G.3 Visualized Analysis
Fig.9 presents a visual comparison among metrics. We show the most error-prone cases for three types of image stitching, where images with green borders represent the correctly stitched images. Conversely, images with blue borders indicate instances where the stitching process resulted in errors or suboptimal output.
The first example is the seam in the stitched image caused by uneven hue. It is evident that Q-Align, Topiq, and UNIQ cannot accurately identify the image with the superior stitching quality in this scenario. According to the design logic of the general NR-IQA models, they prefer the brighter images, leading to the judgement error. While, in our MLLMs-based metrics, prompt can be used to guide the assessment criteria. By incorporating seam-related prompts, our metrics can correctly identify the higher-quality stitched image, thereby overcoming the limitations faced by current NR-IQA models.
The second example is the presence of Artifacts, all evaluated metrics can correctly identify the better stitched image. The appearance of artifacts closely resembles that of ghosting effects, leading to content blurring, which is an important factor in general NR-IQA research. Therefore, NR-IQA method can handle such cases well, and our proposed metrics can also handle such scenarios well.
The third example is a misalignment scene, where it can be seen that the wires are not connected correctly. Unfortunately, all metrics fail to correctly judge better stitched images in this scenario, and UNIQ shows the most significant performance degradation. Despite the broken wires being readily apparent to human observers, it is clear that such an error highly related to image semantics is still a major challenge for existing image quality assessment models. It suggests that future research should explore the integration of more semantically-aware factors into these metrics to improve their efficacy in recognizing and evaluating such defects.
Figure 8: Qualitative results on the different metrics.
Appendix H More Results of Qualitative Evaluations
Due to space limitations, we only present a limited number of qualitative evaluation results in the main paper. Here, to further examine the generalization capabilities of the RDIStitcher, We provide additional qualitative results obtained on the UDIS-D [23], APAPdataset [42], REWdataset [16], and SPWdataset [19].
Fig.9 illustrates the qualitative results for more challenging scenarios. The first example is the large parallax and structuring scene, where the difficulty is how to maintain the original structure during stitching. UDISplus+R shows obvious structural distortion when dealing with such scenes, while SRStitcher shows obvious seams. In contrast, our method achieves a better smoothing effect in the seam region while preserving the object structure; The second example is the structuring and multi-depth layer scene, where the challenge is aligning different depth layers uniformly. Here, it is evident that only the background is correctly aligned, while the foreground runway exhibits misalignment, leading to severe artifacts in UDIS+R. Similarly, both UDISplus+R and SRStitcher show clear misalignments in the white lines of the runway. In contrast, our method excels in this scenario by maintaining the structural integrity of the runway lines; The fourth scene contains some non-rigid transformation objects, which are wires in the sky. These wires do not adhere to rigid transformation principles during the imaging process. As a result, commonly used registration methods, such as homography, face significant challenges in accurately aligning these elements, leading to the problem of abnormal number of wires in UDIS+R and UDISplus+R. In contrast, our method handles this scenario well, demonstrating the robustness of our scheme in dealing with registration errors; The fifth example is also a multi-depth layer scene, UDIS+R struggles with significant artifacts, while UDISplus+R causes severe distortions, particularly affecting the pillars’ structure. In contrast, inpainting-based methods handle this scene effectively, demonstrating their advantage in maintaining structural integrity and reducing distortion. This highlights the superiority of inpainting-based methods in this kind of scene; The sixth example is the scene of large parallax and multi-depth layers, in which UDIS+R still exhibits artifacts, while UDISplus+R is distorted, resulting in structural distortion of the pillars. In contrast, our method demonstrates sustained high performance in such challenging scenarios, effectively maintaining structural accuracy and minimizing distortions.
Fig.10 presents the results under more datasets, including APAPdataset [42], REWdataset [16], and SPWdataset [19].
Figure 9: Additional qualitative results for the challenging scenarios.
Figure 10: Additional qualitative results on APAPdataset [42], REWdataset [16], and SPWdataset [19].